INAF queue observing, 2017 Apr 17-18

Observer: R. Carini, M. Faccini
Telescope Operator: S. Allanson
Support Astronomer: D. Thompson
Instrument: MODS; LUCI

Summary

At the sunset the sky is partly cloudy

Night duration (12\xB0 el twilight) 03:00am-11:50pm UT = 08 h 50m

Observing time: 6 hrs 20 m (72%)

Weather time loss: 0 hrs (0%)

technical time loss 2h 30m (28%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
telluric Marziani F HIP41751 H    
Marziani F SDSSJ0845 H 0.5h completed
telluric Piconcelli C HIP42853 HKspec    
Piconcelli C 0941+3257 HKspec 0.5 completed
Piconcelli D 1013+5615 HKspec 0.5 completed
telluric Piconcelli D,E HIP53735 HKspec    
Piconcelli E 1014 +4300 HKspec 0.3 completed
Grazian F sdss37 dual grating 2h 30m the first two images (1 in mods1 and 1 in mods2) taken with clouds

Details

*Dark

02:36 start the acquisition of telluric for Marziani_F
  • luci1.20170417.0012-0014 , luci2.20170417.0011-0013
  • luci1.20170417.0015-0016 , luci2.20170417.0014-0015 science
  • seeing about 0.7''
  • the time of readout in LUCI1 is negative. we close and open the geirs GUI for LUCI1

02:54 Target F Marziani
  • luci1.20170417.0017 , luci2.20170417.0016
  • the time of readout is negative again, the TO calls Dave Thompson
03:05 the problem now is present also for LUCI2
  • Dave Thompson said us to close the whole LUCI to fix the problem.
  • 03:19 We can open the software again
  • 03:29 we start the alignment, we click reset
  • 03:44 start again the script, jumping the acquisition of the images of the source and the slit
  • during the alignment with LUCI1, the software's frozen. We close it with xkill and open it again

03:58 start again the alignment
  • luci1.0022-0025, luci2.0022-0025

04:09 we stop the alignment, Dave Thompson have to do some checks

04:11 we start again
  • luci1.0027-0031, luci2.0026-0030 science
  • seeing about 0.8''-0.9''
  • now the sky is clear

04:38 slew to the telluric for the target Piconcelli_C
  • luci1.0032-0034, luci2.0031-0033 acquisition
  • luci1.0035-0036, luci2.0034-0035 science

5:00 target C Piconcelli
  • luci1.0037-0040, luci2.0036-0039 acquisition
  • luci1.0041-0043, luci2.0040--0042 science

05:24 target D Piconcelli
  • 05:26 the To checks the pointing
  • 05:30 sent again the script
  • 05:40 LUCI frozen again, we open LUCI in another machine

05:45 we continue the alignment
  • luci1.0044-0047, luci2.0043-0046 acquisition
  • luci1.0048-0050, luci2.0047--0049 science
  • seeing about 1.2''

06:14 killl LUCI in the machine 3 (whre it is frozen)
  • sent the script for the telluric star for Piconcelli D and E
  • luci1.0051-0053, luci2.0050-0052 acquisition
  • luci1.0054-0056, luci2.0053-0055 science

06:30 target E
  • luci1.0057-0059, luci2.0056-0058 acquisition
  • luci1.0060-0061, luci2.0059-0060 science
  • 06:48 seeing about 0.6''

06:56 target F
  • 06:58 check the pointing
  • 07:01 LUCI1 and LUCI2 loses all the connections
  • the TO try to reconnect the instrument and calls Dave Thompson. during the disconnection, the slit is stuck in TURN OUT. it must be put in FPU manually, to avoid the warm up of LUCI
  • we can not switch to MODS until this operation.

08:18 we can switch to MODS

08:38 start Grazian program
  • acqBinoMODS sdss37_ls_newguide.acq * mods1r.20170517.0001-0003 * mods2r.201704017.0001-0004 during the acquisition a lot of clouds come (obviously!!!)

09:10 execBinoMODS sdss37_ls_1.obs * mods1r.20170517.0004 , mods1b.0001 * mods2r.201704017.0005 , mods2b.0001

09:29 error offset pointing!!! we have to sent the preset again, the sky now is clear
  • acqBinoMODS sdss37_ls_newguide.acq
  • seeing 0.6
  • mods1r.20170517.0005-0007
  • mods2r.201704017.0006-0009

09:51 execBinoMODS sdss37_ls_1.obs
  • mods1r.20170517.0008-0011 , mods1b.0002-0005
  • mods2r.201704017.0010-0012 , mods2b.0002-0005
  • seeing 0.6''
  • 10:42 seeing about 0.7''

11:02 slew to the standard star bd33d2642
  • acqBinoMODS bd33d2642_ls.acq
  • mods1r.20170517.0012-0014
  • mods2r.201704017.0013-0015
  • the guide star is wrong (guide star =standard star), we change it
  • we send again the acquisition
  • mods1r.20170517.0015-0016
  • mods2r.201704017.0016-0018

11:28 execBinoMODS bd33d2642_ls.obs
  • mods1r.20170517.0016-0018 , mods1b.0006-0008
  • mods2r.201704017.0018-0020 , mods2b.0006-0008

11:38 slew to Massaro program, target J1719.3 (the target is )
  • acqBinoMODS FGL_J1719.3+1206.acq
  • mods1r.20170517.0019-0020
  • mods2r.201704017.0021-0022
  • we had problems in the identify the target.
  • the sun rises
Topic revision: r9 - 19 Apr 2017, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback